# Evolution and compression in LLMs: On the emergence of human-aligned categorization

### v3

@@ 4.1 English color naming
Figure 2: English color naming experiment with LLMs. (a) IB complexity-accuracy tradeoffs achieved by instruction-tuned LLMs (see Appendix appendix E for all/base models), plotted with respect to the English tradeoff (blue star) and IB theoretical bound (black curve) from Zaslavsky2018Efficient. Models vary widely in their tradeoffs, with larger instruction-tuned models reaching the English point. (b) Color naming systems of English (from Lindsey2014Color) and best-performing LLMs. Each system is shown by its mode map, i.e., it is plotted against the WCS grid (Figure 1a), where each chip is colored by the color-centroid of its modal category. (c) English-alignment (top) and IB complexity (bottom) of all LLMs. Markers are the same as in (a), where a black edge indicates the instruction-tuned model and no edge indicates the base model. Across model families, size and instruction-tuning are associated with higher complexity and better alignment to English.
@@
We begin by eliciting color naming responses from LLMs with English color terms. We then evaluate their IB-efficiency and alignment with the actual English color naming system from Lindsey2014Color. As illustrated in Figure 1a, we consider two variants of this task, a text-only variant where colors are presented as sRGB coordinates, and an image-based variant where colors are presented as a color patch. More details on our procedure can be found in Appendix B.
The resulting systems are shown in Figure 2b, their IB tradeoffs are shown in Figure 2a, and their quantitative evaluation is shown in Figure 2c (see also Figure 7, 8 and 9 in Appendix E). We found that the vast majority of LLMs vary in their complexity and English-alignment, with larger instruction-tuned models achieving better alignment and IB-efficiency. No model aligns perfectly with the English system from Lindsey2014Color, though Gemini-2.0 and Gemma 3 27B (inst.) approximate it closely. Llama 3.3 70B (inst.) and Qwen 2.5 32B (inst.) produce naming systems that are reasonably aligned with English and achieve comparable IB-tradeoffs. We also conducted an analysis on the learning trajectory of Olmo 2 32B (Appendix F), which revealed that English-alignment and complexity increase as a function of learning steps and tokens seen during training. Interestingly, an minimal pair analysis of the multimodal LLMs (Figure 8, Appendix E) revealed that there may be a cutoff of roughly 3 bits of complexity for which processing images rather than text (sRGB coordinates) does not improve alignment.
@@
These findings are somewhat surprising, given that prior work (Abdou2021Can; Patel2022Mappinga; Marjieh2024Large) has shown that language models of a variety of sizes can in principle recover human-aligned representations of color terms. In contrast, we find that many state-of-the-art pretrained LLMs struggle to reproduce any coherent color vocabulary, when presented with colors (encoded as sRGB coordinates) in a constrained naming task in the spirit of the WCS procedure. Although instruction tuning is associated with better performance, it does not guarantee comparable IB-tradeoffs or alignment to English, even for models on the order of $\sim$15B parameters. On the other hand, we were surprised to find that some models–particularly Olmo 2 32B (inst.) (OLMoTeam20252) and Qwen 7 2.5 VL (inst.) (Bai2025Qwen25VL) – produced systems with category structure resembling not English, but instead other, very low-resource languages from the WCS (see Figure 9 in Appendix E). This suggests that although many LLMs struggle to recover the same particular distinctions as English speakers, they may still possess a representation of color aligned with that of humans.
@@
Taken together, these findings indicate that while color categorization in LLMs is limited to very recent, large, instruction-tuned models, from these models there is a clear capacity to develop color categories aligned with human color perception. However, this English color naming task still leaves open a crucial question: is this behavior merely a reflection of imitating patterns in human data, or does it signify a more intrinsic LLM inductive bias towards IB-efficiency in categorization?
@@ 4.2 Iterated in-context language learning (IICLL)
Figure 3: IICLL with LLMs converges to near-optimal IB solutions. The trajectories of Gemini 2.0 Flash (upper left), Gemma 3 27B (upper right), Llama 3.3 70B (lower left) and Qwen 2.5 32B (lower right) are plotted on the information plane (same as Figure 2A), together with the IB tradeoffs across human languages (WCS+English) and human IL data. Small black dots correspond to random initializations of chains with varying number of categories, $k\in\{2,3,4,5,6,14\}$. Thin blue lines correspond to the LLMs’ IICLL trajectories. Gemini captures the complexity range observed across human languages, while the other models converge to lower complexity systems. All models are instruction-tuned.
@@
To investigate whether LLMs possess an efficiency bias that extends beyond learning the specific categories of a language on which they were trained, we turn to our second study, which simulates cultural transmission in LLMs. To this end, we sought to replicate the ILL color naming experiment of Xu2013Cultural as closely as possible, using our IICLL paradign shown in Figure 1b (see Appendix G for more details). We considered only large, instruction tuned models that performed well in the English color naming task for our IICLL experiments: Gemini 2.0, Gemma 3 27B, Qwen 2.5 32B, and Llama 3.3 70B – henceforth, Gemini, Gemma, Qwen and Llama (see Appendix L for an analysis showing that smaller models struggle in IICLL).
@@
Figure 3 shows the resulting trajectories of IICLL chains on the information plane. Gemini develops color naming systems that converge to a similar range of near-optimal IB solutions as the typological patterns of the WCS languages, as well as the final generation systems from the human IL chains from Xu2013Cultural. There is also broad qualitative fit between its final generation IICLL systems and languages from the WCS (see Appendix I for a comparison of mode maps). The other LLMs also develop systems that converge to highly IB-efficient solutions, though they are generally limited to the lower range of complexity observed in the WCS.
@@
Additional quantitative support for these observations is provided in Figure 4. These figures show that over generations, LLM systems become (i) more efficient in their mapping of stimuli to terms, (ii) more similar to naturally occurring human color systems and (iii) more similar to optimal IB systems. Furthermore, LLM IICLL chains converge near the bound relatively quickly (after roughly four generations), parallel to human IL dynamics.
@@
Strikingly, many trajectories from all models initially climb in complexity towards the IB bound before slowly evolving downwards alongside it. This suggests that the capacity to learn and transmit complex yet near-optimally efficient category systems, while strongest for Gemini, is present in all four LLMs. One factor that may drive the difference between Gemini and the other models is that the IICLL task requires very strong in-context learning, as models must integrate dozens of in-context training examples to generalize well. For example, the $k=14$ condition includes $84$ examples, and in this setting most of the LLMs immediately converge to low-complexity solutions. However, given that (1) the terms in our IICLL experiment are nonsense (made up) terms and (2) we give no indication to the model that the stimuli are in fact colors, only that they have “features” (Figure 1c), all four models show an impressive ability to evolve systems (at least transiently) that are aligned with the structure of human color perception.
@@
In order to precisely assess whether the emergent LLM systems were nontrivially efficient and aligned to WCS languages, we conducted a rotation analysis (Regier2007Color) on the final, evolved color systems from our LLM experiments alongside the human data from Xu2013Cultural (Figure 11 in Appendix H). This analysis involved rotating the color term assignments along the hue dimension of the WCS color grid and evaluating the difference in their IB-efficiency and alignment scores. We found that such rotations away from the actual emergent systems lead to a significant decrease in efficiency and alignment scores for Gemini systems. This suggests that Gemini achieves efficient categorization that is not arbitrary, but is constrained by the specific, perceptually grounded organization of human color categories.
Taken together, our IICLL experiments show that LLMs, like human learners, can in principle restructure initially random color naming systems in a process of simulated cultural evolution that leads to near-optimality and high alignment with human semantic representations.
@@ 4.3 The emergence of categories in Shepard circles via IICLL
While many results in color have previously been shown to generalize to other domains (Zaslavsky2019Semantic; Zaslavsky2021Lets; Mollica2021Forms), it is far from trivial to conduct a full-scale analysis of the scope we performed for color naming, primarily due to the lack of high-quality data from both native speakers and ILL experiments. Here, nevertheless, we apply the IICLL paradigm in a qualitatively distinct semantic domain to provide initial evidence that LLMs may have a general bias to learn structured category systems.
We considered a synthetic domain of so-called “Shepard circles,” a classic conceptual space used to study how humans categorize multidimensional stimuli (Shepard1964Attention). These are circles which vary in both radius and angle of rotation of an internal spoke. We generated stimuli by taking eight evenly spaced values for each of the two dimensions, yielding $64$ total stimuli (Figure 5a). For this preliminary investigation, we limit our analysis to Gemini. We found that presenting these stimuli as pairs of numbers (radius and angle) proved to be challenging. This is perhaps unsurprising, because unlike color– which is represented in various ways online in both text and images– there is likely no text online that would allow the model to associate these numbers with meaningful perceptual features. To overcome this limitation and better mimic human perception, we presented Gemini with images of the stimuli.
The results of a sample of IICLL chains are depicted in Figure 5b. Over generations, Gemini transmitted categories that became increasingly compact in their partitioning of the conceptual space, and distinguished regions based on both the radius and angle of the circles. This suggests that LLMs– especially large multimodal models– potentially have a domain-general bias to organize perceptual features into non-arbitrary, and increasingly regular, semantic categories.
@@ 5 Discussion
In this work, we combined a theory-driven approach, based on the IB principle, with cognitively-motivated experimental methods, based on color naming and iterated language learning to study whether LLMs can acquire a human-like inductive bias toward optimally-compressed semantic representations, without being trained for this objective. To do this, we first conducted an in-depth analysis of English color naming across 39 LLMs, and found that a surprising number of state-of-the-art models fail to capture the English color naming system. However, some of the most recent, larger, instruction-tuned models achieve high English-alignment and comparable IB tradeoffs. We then demonstrated that LLMs that do align well with the English color naming system are not merely mimicking patterns in their training data, but rather exhibit a more fundamental capacity to learn human-aligned, efficient color category systems. To do this, we introduced Iterated in-Context Language Learning (IICLL) to simulate cultural transmission of category systems. Over generations of IICLL, LLMs tend to restructure randomly-initialized artificial category systems toward greater IB-efficiency and alignment to human naming systems. We also provide initial evidence that LLMs can develop structured categories over generations of IICLL in a domain distinct from color, suggesting that our results may apply in other semantic domains. Taken together, our current findings suggest that LLMs are capable of evolving perceptually grounded, human-like semantic systems, guided by the same IB-efficiency principle that underlies human languages. Importantly, neither humans nor LLMs are explicitly trained for optimizing the IB objective, suggesting that IB-efficiency may emerge to support intelligent behavior.
@@
Our empirical results open up several important questions for future research. First, while our work demonstrates that cultural transmission alone (via IICLL) can be a sufficient pressure for some LLMs to develop efficient, human-like category systems, a more complete understanding of language evolution requires integrating functional pressure of language use, e.g., via communication. Therefore, an important future direction is to extend our IICLL framework to incorporate communication as a selective pressure, for example, by adopting models that explicitly integrate both transmission and communication (e.g., Kouwenhoven2024Searching). Second, the precise origins of the bias we observe in LLMs toward efficiency are unclear (for example, how might this bias emerge from properties of the training data, instruction-tuning, or model size), and investigating this is another important direction for future work. Lastly, we look forward to extending our analyses across more languages and semantic domains.
Reproducibility statement
@@
This work utilizes a previously published model and code from Zaslavsky2018Efficient, made available at https://github.com/nogazs/ib-color-naming under the MIT License. The ILL data from Xu2013Cultural were obtained from the authors, and the WCS data are publicly available from the World Color Survey website: https://linguistics.berkeley.edu/wcs/. Appendix D provides a list of the specific models used with their Hugging Face or Google API IDs, and Appendix J includes example prompts. We will make the code for our experiments and analyses available upon publication.
@@ Acknowledgments
This work was supported by a Google Cloud Platform Credit Award and a BSF Research Grant (2024318) to NZ. An earlier version of this work was accepted at the NeurIPS 2025 Workshop on Interpreting Cognition in Deep Learning Models (CogInterp).
@@ Appendix A IB Communication Model and Theoretical Bound
The framework is based on a basic communication model that considers an inventory of words $\mathcal{W}$ to communicate about a space of meanings, $\mathcal{M}$. The specific model we apply in this paper is the previously published IB color naming model of Zaslavsky2018Efficient. In this model, meanings are assumed to be mental representations or beliefs over world states $\mathcal{U}$, formally defined as a probability distribution $m(u)$ over world states $u\in\mathcal{U}$. In our setting, the set of world states $\mathcal{U}$ is taken to be the 330 color chips from the WCS grid, and meanings over colors are grounded in the CIELAB perceptual space such that each target color referent, $u_{t}\in\mathcal{U}$, is represented as a Gaussian distribution $m_{t}(u)$ centered around $u_{t}$. Meanings are drawn from an information source, $p(m)$, which characterizes how often each meaning needs to be communicated. The need distribution over meanings in this model, $p(m)$, was estimated using the method of least-informative priors (see Zaslavsky2020Communicative for an extensive evaluation of this need distribution). Note that, in our IICLL experiments, we sample stimulus-word pairs uniformly for training data for each generation, instead of biased sampling from this prior. This was done in order to replicate the procedure in Xu2013Cultural.
Given a meaning $m\sim p(m)$, a speaker produces a signal $w$ using a stochastic production policy, also called an encoder, $q(w|m)$, and then a listener interprets the signal by reconstructing an estimated belief state $\hat{m}_{w}(u)=\sum_{m}q(m|w)m(u)$. Note that this form of listener interpretations corresponds to a Bayesian listener. While this form is assumed here for simplicity, it is not an actual assumption but rather a derivation from the theory (see the SI of Zaslavsky2018Efficient).
@@
A semantic category system in this framework corresponds to a stochastic encoder, $q(w|m)$, which maps meanings to signals. An optimal semantic system, according to the IB principle, is an encoder that satisfies a tradeoff between its informational complexity and communicative accuracy. Complexity, also known as information rate (Cover2006Elements), is defined by the mutual information between speaker meanings and signals,
$I_{q}(M;W)=\sum_{m}p(m)q(w|m)\log\frac{q(w|m)}{q(w)},$ (2)

### v4

@@ 4.1 English color naming
Figure 2: English color naming experiment with LLMs. (a) IB complexity-accuracy tradeoffs achieved by instruction-tuned LLMs (see Appendix E for base models too), plotted in comparison to the English tradeoff (blue star) and the IB bound (black curve) from Zaslavsky et al. (2018). Models vary widely in their tradeoffs, with larger models approaching the English point. (b) Color naming systems of English (Lindsey and Brown, 2014) and best-performing LLMs. Each system is shown by its mode map, i.e., it is plotted against the WCS grid (Figure 1a), where each chip is colored by the color-centroid of its modal category. (c) English-alignment (top) and IB complexity (bottom) of all LLMs. Markers are the same as in (a), where a black edge indicates the instruction-tuned model and no edge indicates the base model. Across model families, size and instruction-tuning are associated with higher complexity and better alignment to English.
@@
We begin with the results from our English color naming study, in which we elicited color naming responses from LLMs with English color terms and then evaluate their IB-efficiency and alignment with the actual English color naming system from Lindsey and Brown (2014). As illustrated in Figure 1a, we consider two variants of this task, a text-only variant where colors are presented as sRGB coordinates, and an image-based variant where colors are presented as a color patch. More details on our procedure can be found in Appendix B.
@@
The resulting systems are shown in Figure 2b, their IB tradeoffs are shown in Figure 2a, and their quantitative evaluation is shown in Figure 2c (see also Figures 7, 8 and 9 in Appendix E). We find that LLMs vary widely in their complexity and English-alignment, with larger instruction-tuned models achieving better alignment and IB-efficiency. No model aligns perfectly with the English system from Lindsey and Brown (2014), but Gemini-2.0 and Gemma 3 27B (inst.) approximate it closely. The fact that so many state-of-the-art pretrained LLMs struggle to recapitulate the English color naming system, even though these models are trained on massive amounts of English data and have billions of parameters, is quite striking. While instruction tuning is generally associated with better performance, it does not guarantee IB-efficiency or alignment to English, even for very large models (for example, Llama 3.3 70B inst.). At the same time, we were surprised to find that some models—particularly Olmo 2 32B (inst.) and Qwen 2.5 VL 7B (inst.)—produced systems with category structure resembling not English, but instead other, very low-resource languages from the WCS (see Figure 9 in Appendix E). This suggests that although many LLMs struggle to recover the same particular distinctions as English speakers, they may still possess human-like color categories.
To better understand how LLMs acquire color categories during training, we analyzed the learning trajectory of Olmo 2 32B (Appendix F). English-alignment only slightly increases during pre-training, and the most substantial improvement occurs in the second, instruction-tuning stage. This finding is consistent with the results of Figure 2c, suggesting that instruction-tuning is an important factor in the models’ ability to exhibit English-like color naming systems.
Finally, to explore the impact of the input representation, we ran two additional analyses. First, we conducted a minimal pair analysis of the multimodal LLMs (Figure 8, Appendix E), comparing systems that were generated with textual vs. image-based representations of color. Interestingly, presenting colors as images rather than textual sRGB coordinates does not improve the alignment or efficiency of larger models, and can even harm their performance, but it does seem to boost the performance of smaller models. Second, we tested the impact of using CIELAB coordinates instead of sRGB, which better captures human perceptual similarities between nearby colors. Consistent with previous findings by Marjieh et al. (2024), we find that all models, including the best performing ones, struggled to align with English naming when colors are presented in CIELAB. This reveals a key difference between how LLMs represent color and how humans do.
@@
Taken together, our findings show that while many LLMs struggle to align with humans on a very simple, but perceptually-grounded, color-naming task, frontier models do exhibit human-aligned color naming systems. This, in turn, opens a crucial question: is this behavior merely a reflection of imitating patterns in the models’ training data, or does it signify a more intrinsic LLM inductive bias towards IB-efficiency in categorization?
@@ 4.2 Iterated in-context language learning (IICLL)
Figure 3: IICLL with LLMs converges to near-optimal IB solutions. The trajectories of Gemini 2.0 (upper left), Gemma 3 27B (upper right), Llama 3.3 70B (lower left) and Qwen 2.5 32B (lower right) are plotted on the information plane (same as Figure 2A), together with the IB tradeoffs across human languages (WCS+English) and human IL data. Small black dots correspond to random initializations of chains with varying number of categories, $k\in\{2,3,4,5,6,14\}$. Thin blue lines correspond to the LLMs’ IICLL trajectories. Gemini captures the complexity range observed across human languages, while the other models converge to lower complexity systems. All models are instruction-tuned.
@@
To investigate whether LLMs possess an efficiency bias that extends beyond learning the specific categories of a language on which they were trained, we turn to our second study, which simulates cultural transmission in LLMs. To this end, we sought to replicate the ILL color naming experiment of Xu et al. (2013) as closely as possible, using our IICLL paradign shown in Figure 1b (see Appendix G for more details). We considered only large, instruction tuned models that performed well in the English color naming task for our IICLL experiments: Gemini 2.0, Gemma 3 27B, Qwen 2.5 32B, and Llama 3.3 70B—henceforth, Gemini, Gemma, Qwen and Llama (see Appendix L for an analysis showing that smaller models struggle in IICLL to produce non-degenerate category systems).
@@
Figure 3 shows the resulting trajectories of IICLL chains on the information plane. Gemini develops color naming systems that converge to a similar range of near-optimal IB solutions as the typological patterns of the WCS languages, as well as the final generation systems from the human IL chains from Xu et al. (2013). There is also broad qualitative fit between its final generation IICLL systems and languages from the WCS (see Appendix I). The other LLMs also develop systems that converge to highly IB-efficient solutions, although they appear to be limited to the lower range of complexity observed in the WCS.
@@
Additional quantitative support for these observations is provided in Figure 4. These figures show that over generations, LLM systems become more efficient in their mapping of stimuli to terms (Figure 4a), more similar to naturally occurring human color systems (Figure 4c), and more similar to optimal IB systems (Figure 4b). Furthermore, LLM IICLL chains converge near the bound relatively quickly (after roughly four generations), parallel to human IL dynamics.
@@
Strikingly, many trajectories from all models initially climb in complexity towards the IB bound before slowly evolving downwards alongside it. This suggests that the capacity to learn and transmit complex yet near-optimally efficient category systems, while strongest for Gemini, is present in all four LLMs. One factor that may drive the difference between Gemini and the other models is that the IICLL task requires very strong in-context learning, as models must integrate dozens of in-context training examples to generalize well. For example, the $k=14$ condition includes $84$ examples, and in this setting most of the LLMs immediately converge to low-complexity solutions. However, given that the terms in our IICLL experiment are made up (pseudo) terms and that we give no indication to the model that the stimuli are in fact colors, only that they have “features” (Figure 1c), all four models show an impressive ability to evolve IB-efficient systems.
@@
To assess whether the emergent LLM systems are non-trivially efficient and aligned to WCS languages, we conducted a rotation analysis (Regier et al., 2007) on the final, evolved color systems from our LLM experiments alongside the human data from Xu et al. (2013) (Figure 11 in Appendix H). This analysis involved rotating the color-label mapping along the hue dimension of the WCS color grid (i.e., along the columns of Figure 1a) and evaluating the difference in their efficiency and alignment scores. We find that such rotations away from the actual emergent systems lead to a significant decrease in efficiency and alignment for Gemini, while the results are less conclusive for the other models. Furthermore, we find that Gemini’s efficiency and alignment over generations are higher than the human IL trajectories, and are also higher than a baseline learner based on an alternative feature-based clustering algorithm (Appendix M). This suggests that Gemini truly exhibits an emergent inductive learning bias toward IB-efficiency, even though as fas as we know it was not trained or fine-tuned with respect to the IB objective function.
@@ 4.3 The emergence of categories in Shepard circles via IICLL
While many results in color have previously been shown to generalize to other domains (Zaslavsky et al., 2019; 2021; Mollica et al., 2021), it is challenging to conduct in other domains a full-scale analysis of the scope we performed for color naming, primarily due to the lack of high-quality data from both native speakers and ILL experiments. Here, nevertheless, we apply the IICLL paradigm in a qualitatively distinct semantic domain to provide initial evidence that our results may indeed generalize beyond color.
To this end, we considered a synthetic domain of so-called “Shepard circles,” a classic conceptual space used to study how humans categorize multidimensional stimuli (Shepard, 1964). These are circles which vary in both radius and angle of rotation of an internal spoke. Following Carr et al. (2020), we generated stimuli by taking eight evenly spaced values for each of the two dimensions, yielding $64$ total stimuli (Figure 5a). For this preliminary investigation, we limit our analysis to Gemini and $k=4$ allowed labels. We found that presenting these stimuli as pairs of numbers (radius and angle) proved to be challenging. This is perhaps unsurprising, because unlike color—which is frequently represented in various ways online in both text and images—there is likely no text online that would allow the model to associate these numbers with meaningful perceptual features. To overcome this limitation and better mimic human perception, we presented Gemini with images of the stimuli. Four samples of IICLL chains are are shown in Figure 5b. Over generations, Gemini transmitted categories that became increasingly compact in their partitioning of the space, and distinguished regions based on both the radius and angle of the circles. This suggests that LLMs—especially frontier multimodal models—potentially have a domain-general bias to organize features into non-arbitrary, and increasingly regular, semantic categories. An important direction for future work is to test whether this emergent structure also supports greater IB-efficiency as seen in humans (Imel et al., 2025).
@@ 5 Discussion
In this work, we combined a theory-driven approach, based on the IB principle, with cognitively-motivated experimental methods, based on color naming and iterated language learning to study whether LLMs can acquire a human-like inductive bias toward optimally-compressed semantic representations, without being trained for this objective. To do this, we first conducted an in-depth analysis of English color naming across 39 LLMs, and found that a surprising number of state-of-the-art models fail to capture the English color naming system. However, some of the most recent, larger, instruction-tuned models achieve high English-alignment and comparable IB tradeoffs. We then demonstrated that LLMs that do align well with the English color naming system are not merely mimicking patterns in their training data, but rather exhibit a more fundamental capacity to learn human-aligned, efficient color category systems. To do this, we introduced Iterated in-Context Language Learning (IICLL) to simulate cultural transmission of category systems. Over generations of IICLL, LLMs tend to restructure randomly-initialized artificial category systems toward greater IB-efficiency and alignment to human naming systems. We also provide initial evidence that LLMs can develop structured categories over generations of IICLL in a domain distinct from color, suggesting that our results may apply in other semantic domains as well. Taken together, our findings suggest that LLMs are capable of evolving perceptually grounded, human-like semantic systems, guided by the same IB-efficiency principle that underlies human languages. Importantly, neither humans nor LLMs are explicitly trained for optimizing the IB objective, suggesting that IB-efficiency may emerge to support intelligent behavior.
@@
Our empirical results open up several important questions for future research. First, while our work demonstrates that cultural transmission alone (via IICLL) can be a sufficient pressure for some LLMs to develop efficient, human-like category systems, a more complete understanding of language evolution requires integrating functional pressure of language use, e.g., via communication. Therefore, an important future direction is to extend our IICLL framework to incorporate communication as a selective pressure, for example, by adopting models that explicitly integrate both transmission and communication (e.g., Kouwenhoven et al., 2024). Second, the precise origins of the bias we observe in LLMs toward efficiency are unclear (for example, how might this bias emerge from properties of the training data, instruction-tuning, or model size), and investigating this is another important direction for future work. Lastly, it is important to extend our analyses across more languages and semantic domains.
Reproducibility statement.
@@
This work utilizes a previously published model and code from Zaslavsky et al. (2018), made available at https://github.com/nogazs/ib-color-naming under the MIT License, and code from Carr et al. (2020), made available from https://github.com/jwcarr/shepard under a CC BY 4.0 License. The ILL data from Xu et al. (2013) were obtained from the authors, and the WCS data are publicly available from the World Color Survey website, https://linguistics.berkeley.edu/wcs. Appendix D provides a list of the specific models used with their Hugging Face or Google API IDs, and Appendix J includes example prompts. Code for reproducing our experiments can be found at https://infocoglab.github.io/evolution-compression-llms. To support full reproducibility, the LLM data that was generated for this paper is available upon request.
@@ Acknowledgments
This work was supported by a Google Cloud Platform Credit Award and a BSF Research Grant (2024318) to NZ. An earlier version of this work received the best paper award at the NeurIPS 2025 Workshop on Interpreting Cognition in Deep Learning Models (CogInterp).
@@ Appendix A IB Communication Model and Theoretical Bound
The IB framework is based on a basic communication model that considers an inventory of words $\mathcal{W}$ to communicate about a space of meanings, $\mathcal{M}$. The specific model we apply in this paper is the previously published IB color naming model of Zaslavsky et al. (2018). In this model, meanings are assumed to be mental representations or beliefs over world states $\mathcal{U}$, formally defined as probability distributions $m(u)$ over world states $u\in\mathcal{U}$. In our setting, the set of world states $\mathcal{U}$ is taken to be the 330 color chips from the WCS grid, and meanings over colors are grounded in the CIELAB perceptual space such that each target color referent, $u_{t}\in\mathcal{U}$, is represented as a Gaussian distribution $m_{t}(u)$ centered around $u_{t}$. Meanings are drawn from an information source, $p(m)$, which characterizes how often each meaning needs to be communicated. The need distribution over meanings in this model, $p(m)$, was estimated using the method of least-informative priors (see Zaslavsky et al. (2020) for an extensive evaluation of this need distribution). Note that, in our IICLL experiments, we sample stimulus-word pairs uniformly for training data for each generation, instead of biased sampling from this prior. This was done in order to replicate the procedure in Xu et al. (2013).
Given a meaning $m\sim p(m)$, a speaker produces a signal $w$ using a stochastic production policy, also called an encoder, $q(w|m)$, and then a listener interprets the signal by reconstructing an estimated belief state $\hat{m}_{w}(u)=\sum_{m}q(m|w)m(u)$. Note that this form of listener interpretations corresponds to a Bayesian listener. While this form is assumed here for simplicity, it is not an actual assumption but rather a derivation from the theory (see the SI of Zaslavsky et al. (2018)).
@@
A semantic category system in this framework corresponds to a stochastic encoder, $q(w|m)$, which maps meanings to signals. An optimal semantic system, according to the IB principle, is an encoder that satisfies a tradeoff between its informational complexity and communicative accuracy. Complexity, also known as information rate (Cover and Thomas, 2006), is defined by the mutual information between speaker meanings and signals,
$I_{q}(M;W)=\sum_{m,w}p(m)q(w|m)\log\frac{q(w|m)}{q(w)},$ (2)